OpenStudy (anonymous):

A sum of money invested at 3% per 6-month period will double in approximately how many years?

OpenStudy (amistre64):

you essentially have to log it out ..

OpenStudy (amistre64):

is the 3% a period rate? it reads as tho its 3% every 6 months .. how many periods does it take? 2P = P(1.03)^n 2 = (1.03)^n log(2) = n log(1.03) etc ...

OpenStudy (anonymous):

Well, the question is given just like that. But I'll try log. Thanks.

OpenStudy (amistre64):

logs allow us to isolate an exponent ...

OpenStudy (amistre64):

23.4 is the number of periods, each period is 1/2 a year

OpenStudy (anonymous):

Gotcha. 23.4/2= 11.7

OpenStudy (amistre64):

we have a 2:1 ratio ... thats it

OpenStudy (anonymous):

Thank you.

OpenStudy (amistre64):

youre welcome